# Derivation of variational membrane models in the context of anisotropic nonlocal hyperelasticity

### v1

@@ 1. Introduction
where $u$ describes the deformation of the reference configuration $\Omega\subset{\mathbb{R}}^{n}$, and $W\colon{\mathbb{R}}^{n\times n}\to[0,\infty)$ denotes the material energy density. These models were already proposed in [58, Section 18], but have only recently been subjected to a rigorous variational analysis. The first work to investigate the Sobolev spaces associated with the nonlocal gradients was [56] (see also [57, 21]), in which the Riesz fractional gradient was considered that arises by choosing the interaction kernel $\rho=|\cdot|^{-(n+s-1)}$ for some fractional exponent $s\in(0,1)$. There, the authors established fractional Poincaré inequalities and compact embeddings, and addressed existence of minimizers for convex energy densities. These existence results were later generalized to polyconvex and quasiconvex energy densities [5, 37], which are typically employed in applications of hyperelasticity. To better align with peridynamic models, the extension to finite-horizon fractional gradients was considered in [8, 24]. More recently, a theory for general compactly supported interaction kernels $\rho$ was developed [11], a setting that we also adopt in this paper. Beyond the results on the function spaces and the existence of minimizers, it has furthermore been shown in [25] (see also [45, 10]) that nonlocal models $\Gamma$-converge to local ones as the horizon vanishes, in particular for families of kernels $\rho_{\delta}:=\delta^{-n}\rho(\cdot/\delta)$ for $\delta>0$.
@@
We are particularly interested in the study of thin structures, which are known to show fundamentally different responses to external loading compared to bulk materials. A natural way to analytically capture these effects within classical nonlinear hyperelasticity is to study the asymptotic behavior of three-dimensional energy functionals on domains whose thickness tends to zero and to identify the resulting lower-dimensional models. Depending on the underlying scaling, one obtains qualitatively different theories in the limit. Early rigorous results in this direction include limits that yield string [1] and membrane models [42]. The subsequent analysis of thin structures was strongly shaped by the seminal paper [33], which led, among other things, to a systematic derivation of plate theories in several scaling regimes [33, 34] and the generalization to non-Euclidean elasticity [43], as well as the development of rod theories [46, 47, 52, 53]. In the last two decades, variational dimension reduction has seen extensions in several different directions. A recurring theme is the incorporation of non-convex differential constraints, motivated for instance by incompressibility [22, 23, 30, 31], non-self-interpenetration of matter and (global) invertibility [41, 49, 54, 12], or $\mathcal{A}$}-free fields [36, 39] with applications in elasticity or micromagnetism. As the latter indicates, thin-structure limits have also been pursued well beyond the purely elastic setting. Further extensions include elastoplasticity [32, 26, 27], magnetoelasticity [17, 16], liquid crystal elastomer sheets [63], or fracture mechanics [3, 55, 2], to name just a few.
@@
A technical and modeling nuisance for the energies (1.3) and (1.4) is their invariance under the addition of functions with vanishing gradient. In the setting $\delta=0$, this invariance is harmless, since the only functions with zero gradient are constants. The nonlocal setting, however, gives rise to highly nontrivial functions $h\in H^{\rho,p}_{\delta^{\varepsilon}}(\Omega;{\mathbb{R}}^{3})$ with $D_{\rho,\delta^{\varepsilon}}h=0$, see Remark 2.6 and [38]; as a consequence, such large deformations are energetically invisible. In the spirit of [60, Section 6], we therefore add a stabilization term that penalizes such contributions and, in particular, discourages them from appearing along minimizing, or almost-minimizing, sequences:
@@ 2. Anisotropic nonlocal gradients
b) Suppose $O$ is a bounded Lipschitz domain and $\delta=(1,1)$. Then, it holds that $N^{\rho,p}(O;{\mathbb{R}}^{3}):=N^{\rho,p}_{\delta}(O;{\mathbb{R}}^{3})$ is an infinite-dimensional space and hence, contains many non-constant functions. Indeed, this can be argued similarly as in the case for truncated fractional kernels from [38, Proposition 3.3] using the more general results from [11]. In fact, in the case that $\rho$ is as in Example 2.1 with $\chi$ locally constant around the origin, $p\in(1,\frac{2}{1-s})$ and $O$ is a $C^{1,1}$-domain, then it follows from [38, Theorem 3.8] that for every boundary condition $g\in L^{p}(O_{\delta}\setminus O;{\mathbb{R}}^{3})$ and every $c\in{\mathbb{R}}^{3}$ there is a unique $h\in N^{\rho,p}(O;{\mathbb{R}}^{3})$ with $h=g$ in $O_{\delta}\setminus O$ and
@@
This characterizes the set $N^{\rho,p}(O;{\mathbb{R}}^{3})$ completely. When $p\geq\frac{2}{1-s}$, uniqueness still holds but existence may fail due to possible singularities forming at the boundary $\partial O$. This characterization does not readily extend to general kernels $\rho$, but we mention at least that uniqueness for $p\geq 2$ is still true by following the argument in [38, Proposition 3.9] together with [11, Lemma 4.3]. These facts will not be needed for our analysis though.
@@
We now deal with the Leibniz rule of smooth functions and ones in $H^{\rho,p}_{\delta}(\Omega;{\mathbb{R}}^{3})$ for open and bounded sets $\Omega\subset{\mathbb{R}}^{3}$. The proof essentially follows the structure of [24, Lemma 12] and uses the calculations and estimates in Lemma 2.9.
@@
Having established the Leibniz rule, we turn to the density of $C_{c}^{\infty}({\mathbb{R}}^{3};{\mathbb{R}}^{3})$ in $H^{\rho,p}_{\delta}(\Omega;{\mathbb{R}}^{3})$ on cylindrical bounded Lipschitz domains $\Omega$. After a regularization in the out-of-plane variable, the proof follows the same arguments as in [24, Theorem 1 and Lemma 13].
@@
The case $\delta\in(0,1]^{2}$ can be argued as in [24, Theorem 1] using the Leibniz rule in Lemma 2.9 and 2.11 a) and does not even require the Cartesian product structure of $\Omega$. The case $\delta=(0,0)$ is the classical density result for Sobolev spaces. It remains to prove the case $\bar{\delta}>0$ and $\delta_{3}=0$ and the case $\bar{\delta}=0$ and $\delta_{3}>0$. Since the proof strategy is analogous for both cases, we only detail the former case and split the proof into two steps.
@@
Step 2: Smoothing the remaining variables. If $\chi\in C_{c}^{\infty}({\mathbb{R}}^{2})$, then we can use Lemma 2.11 b) ii)ii) to conclude that $\chi u_{k}\colon x\mapsto\chi(\bar{x})u_{k}(x)$ satisfies $\chi u_{k}\in H^{\rho,p}_{\delta}(\omega^{\prime}\times(0,1);{\mathbb{R}}^{3})$ as long as $(\omega^{\prime}\setminus\omega)\cap\operatorname{supp}(\chi)=\emptyset$; here, we crucially used that the third partial derivative of $u_{k}$ exists in the weak sense. Hence, we can completely follow the proof of [24, Theorem 1 and Lemma 13] using partition of unity, translation and mollification – albeit only in the first two variables – in order to be able to approximate each $u_{k}$ with a sequence of functions $(u_{k,j})_{j}$ whose weak in-plane gradients exist of all orders and are in $L^{p}$. Combining this with the fact that the weak derivative $\partial_{3}^{n}u_{k}$ exists for all $n\in{\mathbb{N}}$ yields that $u_{k,j}\in W^{n,p}({\mathbb{R}}^{3};{\mathbb{R}}^{3})$ for all $n\in{\mathbb{N}}$ and all $u_{k,j}$ vanish outside of a compact set (uniformly in $k,j$). Standard Sobolev embedding results yield that all of these functions are actually smooth. A diagonalization argument then finishes the proof. ∎
@@
which is the inverse of ${\mathcal{Q}}_{\rho}:={\mathcal{Q}}_{\rho,(1,1)}$, see [25, Lemma 2.12]. Moreover, for $\varphi\in C_{c}^{\infty}({\mathbb{R}}^{3};{\mathbb{R}}^{3})$, it holds by [19, Equation (31)] that
@@
and $\operatorname{div}V_{\rho}$ is a radial function which agrees with a Schwartz function on $B_{1}(0)^{c}$, see [19, Lemma 4.8]. We claim that
@@
where the first and third inequality use a bound on translations in (fractional) Sobolev spaces, see e.g. [18, Proposition B.2], and the second inequality is the interpolation inequality (1.7) for $s=\lambda_{2}$, cf. [9, Corollary 3.19]. We now obtain as before with Minkowski’s integral inequality that
@@
(iii) This can be argued as in [38, Theorem 4.1] by using Part (i) and (ii). ∎
@@
This is an immediate consequence of Proposition 2.15 (iii), cf. [38, Lemma 4.7] for more details. ∎
@@ 3. Thin films in nonlocal hyperelasticity
Step 2: Liminf-inequality. Given the previous compactness step, we now define $v_{\varepsilon}:={\mathcal{Q}}_{\rho,\delta^{\varepsilon}}(u_{\varepsilon}-h_{\varepsilon})$ for $\varepsilon>0$ and $v:={\mathcal{Q}}_{\rho,\delta^{0}}u$ with ${\mathcal{Q}}_{\rho,\delta^{\varepsilon}}$ and ${\mathcal{Q}}_{\rho,\delta^{0}}$ as in Proposition 2.15. Then, it holds that $\nabla v_{\varepsilon}=D_{\rho,\delta^{\varepsilon}}u_{\varepsilon}$ and $\nabla v=D_{\rho,\delta^{0}}u$. We then use the bound in Proposition 2.15 (i) and Lemma 2.18 to infer that $v_{\varepsilon}\rightharpoonup v$ in $W^{1,p}(\Omega;{\mathbb{R}}^{3})$. In the spirit of the classical dimension reduction result from [42, Theorem 2], we now deduce that
@@
Step 3: Recovery sequence. Let $u\in H^{\rho,p}_{\delta^{0}}(\Omega;{\mathbb{R}}^{3})$ with $(D_{\rho,\delta^{0}}u)e_{3}=0$ in $\Omega$, then we can define $v:={\mathcal{Q}}_{\rho,\delta^{0}}u\in W^{1,p}(\Omega;{\mathbb{R}}^{3})$ such that $\partial_{3}v=0$. Using the classical recovery sequence from [42, Theorem 2] (without boundary conditions), we find a bounded sequence $(v_{\varepsilon})_{\varepsilon}\subset W^{1,p}(\Omega;{\mathbb{R}}^{3})$ such that $v_{\varepsilon}\to v$ in $L^{p}(\Omega;{\mathbb{R}}^{3})$ and
@@
While this is a mathematically satisfactory result, from a practical perspective, having to deal with the functions in $N^{\rho,p}_{\delta^{\varepsilon}}(\Omega;{\mathbb{R}}^{3})$ is quite unnatural; not least for the fact that this implies that there are non-constant deformations which do not incur additional stored energy. Therefore, we will now also pursue an approach where we include a penalty or regularization term in the spirit of [60, Section 6], which accounts for the functions in $N^{\rho,p}_{\delta^{\varepsilon}}(\Omega;{\mathbb{R}}^{3})$. This should stabilize the functional so that the compactness and convergence of minimizers holds without taking different representatives.

### v2

@@ 1. Introduction
where $u$ describes the deformation of the reference configuration $\Omega\subset{\mathbb{R}}^{n}$, and $W\colon{\mathbb{R}}^{n\times n}\to[0,\infty)$ denotes the material energy density. These models were already proposed in [60, Section 18], but have only recently been subjected to a rigorous variational analysis. The first work to investigate the Sobolev spaces associated with the nonlocal gradients was [58] (see also [59, 23]), in which the Riesz fractional gradient was considered that arises by choosing the interaction kernel $\rho=|\cdot|^{-(n+s-1)}$ for some fractional exponent $s\in(0,1)$. There, the authors established fractional Poincaré inequalities and compact embeddings, and addressed existence of minimizers for convex energy densities. These existence results were later generalized to polyconvex and quasiconvex energy densities [5, 39], which are typically employed in applications of hyperelasticity. To better align with peridynamic models, the extension to finite-horizon fractional gradients was considered in [8, 26]. More recently, a theory for general compactly supported interaction kernels $\rho$ was developed [11], a setting that we also adopt in this paper. Beyond the results on the function spaces and the existence of minimizers, it has furthermore been shown in [27] (see also [47, 10]) that nonlocal models $\Gamma$-converge to local ones as the horizon vanishes, in particular for families of kernels $\rho_{\delta}:=\delta^{-n}\rho(\cdot/\delta)$ for $\delta>0$.
@@
We are particularly interested in the study of thin structures, which are known to show fundamentally different responses to external loading compared to bulk materials. A natural way to analytically capture these effects within classical nonlinear hyperelasticity is to study the asymptotic behavior of three-dimensional energy functionals on domains whose thickness tends to zero and to identify the resulting lower-dimensional models. Depending on the underlying scaling, one obtains qualitatively different theories in the limit. Early rigorous results in this direction include limits that yield string [1] and membrane models [44]. The subsequent analysis of thin structures was strongly shaped by the seminal paper [35], which led, among other things, to a systematic derivation of plate theories in several scaling regimes [35, 36] and the generalization to non-Euclidean elasticity [45], as well as the development of rod theories [48, 49, 54, 55]. In the last two decades, variational dimension reduction has seen extensions in several different directions. A recurring theme is the incorporation of non-convex differential constraints, motivated for instance by incompressibility [24, 25, 32, 33], non-self-interpenetration of matter and (global) invertibility [43, 51, 56, 12], or $\mathcal{A}$}-free fields [38, 41] with applications in elasticity or micromagnetism. As the latter indicates, thin-structure limits have also been pursued well beyond the purely elastic setting. Further extensions include elastoplasticity [34, 28, 29], magnetoelasticity [18, 17], liquid crystal elastomers [22, 13], or fracture mechanics [3, 57, 2], to name just a few.
@@
A technical and modeling nuisance for the energies (1.3) and (1.4) is their invariance under the addition of functions with vanishing gradient. In the setting $\delta=0$, this invariance is harmless, since the only functions with zero gradient are constants. The nonlocal setting, however, gives rise to highly nontrivial functions $h\in H^{\rho,p}_{\delta^{\varepsilon}}(\Omega;{\mathbb{R}}^{3})$ with $D_{\rho,\delta^{\varepsilon}}h=0$, see Remark 2.6 and [40]; as a consequence, such large deformations are energetically invisible. In the spirit of [62, Section 6], we therefore add a stabilization term that penalizes such contributions and, in particular, discourages them from appearing along minimizing, or almost-minimizing, sequences:
@@ 2. Anisotropic nonlocal gradients
b) Suppose $O$ is a bounded Lipschitz domain and $\delta=(1,1)$. Then, it holds that $N^{\rho,p}(O;{\mathbb{R}}^{3}):=N^{\rho,p}_{\delta}(O;{\mathbb{R}}^{3})$ is an infinite-dimensional space and hence, contains many non-constant functions. Indeed, this can be argued similarly as in the case for truncated fractional kernels from [40, Proposition 3.3] using the more general results from [11]. In fact, in the case that $\rho$ is as in Example 2.1 with $\chi$ locally constant around the origin, $p\in(1,\frac{2}{1-s})$ and $O$ is a $C^{1,1}$-domain, then it follows from [40, Theorem 3.8] that for every boundary condition $g\in L^{p}(O_{\delta}\setminus O;{\mathbb{R}}^{3})$ and every $c\in{\mathbb{R}}^{3}$ there is a unique $h\in N^{\rho,p}(O;{\mathbb{R}}^{3})$ with $h=g$ in $O_{\delta}\setminus O$ and
@@
This characterizes the set $N^{\rho,p}(O;{\mathbb{R}}^{3})$ completely. When $p\geq\frac{2}{1-s}$, uniqueness still holds but existence may fail due to possible singularities forming at the boundary $\partial O$. This characterization does not readily extend to general kernels $\rho$, but we mention at least that uniqueness for $p\geq 2$ is still true by following the argument in [40, Proposition 3.9] together with [11, Lemma 4.3]. These facts will not be needed for our analysis though.
@@
We now deal with the Leibniz rule of smooth functions and ones in $H^{\rho,p}_{\delta}(\Omega;{\mathbb{R}}^{3})$ for open and bounded sets $\Omega\subset{\mathbb{R}}^{3}$. The proof essentially follows the structure of [26, Lemma 12] and uses the calculations and estimates in Lemma 2.9.
@@
Having established the Leibniz rule, we turn to the density of $C_{c}^{\infty}({\mathbb{R}}^{3};{\mathbb{R}}^{3})$ in $H^{\rho,p}_{\delta}(\Omega;{\mathbb{R}}^{3})$ on cylindrical bounded Lipschitz domains $\Omega$. After a regularization in the out-of-plane variable, the proof follows the same arguments as in [26, Theorem 1 and Lemma 13].
@@
The case $\delta\in(0,1]^{2}$ can be argued as in [26, Theorem 1] using the Leibniz rule in Lemma 2.9 and 2.11 a) and does not even require the Cartesian product structure of $\Omega$. The case $\delta=(0,0)$ is the classical density result for Sobolev spaces. It remains to prove the case $\bar{\delta}>0$ and $\delta_{3}=0$ and the case $\bar{\delta}=0$ and $\delta_{3}>0$. Since the proof strategy is analogous for both cases, we only detail the former case and split the proof into two steps.
@@
Step 2: Smoothing the remaining variables. If $\chi\in C_{c}^{\infty}({\mathbb{R}}^{2})$, then we can use Lemma 2.11 b) ii)ii) to conclude that $\chi u_{k}\colon x\mapsto\chi(\bar{x})u_{k}(x)$ satisfies $\chi u_{k}\in H^{\rho,p}_{\delta}(\omega^{\prime}\times(0,1);{\mathbb{R}}^{3})$ as long as $(\omega^{\prime}\setminus\omega)\cap\operatorname{supp}(\chi)=\emptyset$; here, we crucially used that the third partial derivative of $u_{k}$ exists in the weak sense. Hence, we can completely follow the proof of [26, Theorem 1 and Lemma 13] using partition of unity, translation and mollification – albeit only in the first two variables – in order to be able to approximate each $u_{k}$ with a sequence of functions $(u_{k,j})_{j}$ whose weak in-plane gradients exist of all orders and are in $L^{p}$. Combining this with the fact that the weak derivative $\partial_{3}^{n}u_{k}$ exists for all $n\in{\mathbb{N}}$ yields that $u_{k,j}\in W^{n,p}({\mathbb{R}}^{3};{\mathbb{R}}^{3})$ for all $n\in{\mathbb{N}}$ and all $u_{k,j}$ vanish outside of a compact set (uniformly in $k,j$). Standard Sobolev embedding results yield that all of these functions are actually smooth. A diagonalization argument then finishes the proof. ∎
@@
which is the inverse of ${\mathcal{Q}}_{\rho}:={\mathcal{Q}}_{\rho,(1,1)}$, see [27, Lemma 2.12]. Moreover, for $\varphi\in C_{c}^{\infty}({\mathbb{R}}^{3};{\mathbb{R}}^{3})$, it holds by [20, Equation (31)] that
@@
and $\operatorname{div}V_{\rho}$ is a radial function which agrees with a Schwartz function on $B_{1}(0)^{c}$, see [20, Lemma 4.8]. We claim that
@@
where the first and third inequality use a bound on translations in (fractional) Sobolev spaces, see e.g. [19, Proposition B.2], and the second inequality is the interpolation inequality (1.7) for $s=\lambda_{2}$, cf. [9, Corollary 3.19]. We now obtain as before with Minkowski’s integral inequality that
@@
(iii) This can be argued as in [40, Theorem 4.1] by using Part (i) and (ii). ∎
@@
This is an immediate consequence of Proposition 2.15 (iii), cf. [40, Lemma 4.7] for more details. ∎
@@ 3. Thin films in nonlocal hyperelasticity
Step 2: Liminf-inequality. Given the previous compactness step, we now define $v_{\varepsilon}:={\mathcal{Q}}_{\rho,\delta^{\varepsilon}}(u_{\varepsilon}-h_{\varepsilon})$ for $\varepsilon>0$ and $v:={\mathcal{Q}}_{\rho,\delta^{0}}u$ with ${\mathcal{Q}}_{\rho,\delta^{\varepsilon}}$ and ${\mathcal{Q}}_{\rho,\delta^{0}}$ as in Proposition 2.15. Then, it holds that $\nabla v_{\varepsilon}=D_{\rho,\delta^{\varepsilon}}u_{\varepsilon}$ and $\nabla v=D_{\rho,\delta^{0}}u$. We then use the bound in Proposition 2.15 (i) and Lemma 2.18 to infer that $v_{\varepsilon}\rightharpoonup v$ in $W^{1,p}(\Omega;{\mathbb{R}}^{3})$. In the spirit of the classical dimension reduction result from [44, Theorem 2], we now deduce that
@@
Step 3: Recovery sequence. Let $u\in H^{\rho,p}_{\delta^{0}}(\Omega;{\mathbb{R}}^{3})$ with $(D_{\rho,\delta^{0}}u)e_{3}=0$ in $\Omega$, then we can define $v:={\mathcal{Q}}_{\rho,\delta^{0}}u\in W^{1,p}(\Omega;{\mathbb{R}}^{3})$ such that $\partial_{3}v=0$. Using the classical recovery sequence from [44, Theorem 2] (without boundary conditions), we find a bounded sequence $(v_{\varepsilon})_{\varepsilon}\subset W^{1,p}(\Omega;{\mathbb{R}}^{3})$ such that $v_{\varepsilon}\to v$ in $L^{p}(\Omega;{\mathbb{R}}^{3})$ and
@@
While this is a mathematically satisfactory result, from a practical perspective, having to deal with the functions in $N^{\rho,p}_{\delta^{\varepsilon}}(\Omega;{\mathbb{R}}^{3})$ is quite unnatural; not least for the fact that this implies that there are non-constant deformations which do not incur additional stored energy. Therefore, we will now also pursue an approach where we include a penalty or regularization term in the spirit of [62, Section 6], which accounts for the functions in $N^{\rho,p}_{\delta^{\varepsilon}}(\Omega;{\mathbb{R}}^{3})$. This should stabilize the functional so that the compactness and convergence of minimizers holds without taking different representatives.